# The Map Equation Goes Neural: Mapping Network Flows with Graph Neural Networks

### v3

@@ Abstract
Community detection is an essential tool for unsupervised data exploration and revealing the organisational structure of networked systems. With a long history in network science, community detection typically relies on objective functions, optimised with custom-tailored search algorithms, but often without leveraging recent advances in deep learning. Recently, first works have started incorporating such objectives into loss functions for neural graph clustering and pooling. We consider the map equation, a popular information-theoretic objective function for unsupervised community detection, and express it in differentiable tensor form for optimisation through gradient descent. Our formulation turns the map equation compatible with any neural network architecture, enables end-to-end learning, incorporates node features, and chooses the optimal number of clusters automatically, all without requiring explicit regularisation. Applied to unsupervised graph clustering tasks, we achieve competitive performance against state-of-the-art neural graph clustering baselines in synthetic and real-world datasets.
@@ 1 Introduction
Many real-world networked systems are organised in communities: groups of nodes that are more similar to each other than to the rest. Communities provide insights into network structure at the mesoscale, revealing sub-systems by analysing link patterns. Motivated by different research questions, several characterisations of what constitutes “good” communities have been proposed [1, 2], however, neither of them is fundamentally more correct than any other. Moreover, no single community-detection method outperforms all others on any given network [3], motivating the ongoing efforts of research on community detection. Typically, community-detection approaches formulate an objective function that calculates a quality score for a given partition of the network’s nodes into communities. Finding the best partition is an NP-hard search problem and often involves custom heuristic algorithms that attempt to minimise their objective function.
@@
Here, we combine the benefits of traditional community-detection approaches and deep learning and consider the map equation, an information-theoretic objective function for community detection [13]. By adapting the map equation for soft cluster assignments and implementing it in differentiable tensor form, we enable end-to-end optimisation of the map equation as a loss function with gradient descent and GNNs. In analogy to the map equation’s stochastic optimisation algorithm Infomap [14], we call our approach Neuromap and evaluate it against Infomap and several recent GNN-based graph clustering methods. Applied to synthetic and real-world networks, Neuromap demonstrates competitive performance against recent neural graph clustering baselines.
@@
    We adapt the map equation as a differentiable loss function for end-to-end neural graph clustering and propose Neuromap, a deep-learning-based alternative to the popular Infomap algorithm for unsupervised community detection with the map equation. Neuromap is compatible with any neural network architecture, detects overlapping communities, leverages node features for improved performance on real-world networks, and, by following the minimum description length principle, does not require explicit regularisation.
@@
    We extensively evaluate Neuromap on hundreds of synthetic and ten real datasets against recent baselines paired with various neural network architectures. Neuromap outperforms the baseline on the synthetic networks and performs best in seven out of ten real datasets.
@@
    By choosing a higher maximum number of clusters than previous works, we show empirically that recent baselines tend to overfit and report considerably more than the ground-truth number of communities. Moreover, we find that choosing a small maximum number of communities is detrimental to graph clustering performance.
@@ 3 Background: the map equation
Let $G=\left(V,E,\delta\right)$ be a graph with nodes $V$, links $E$, and $\delta\colon E\to\mathbb{R}_{0}^{+}$ a function that specifies non-negative link weights. When all nodes are assigned to the same module, the codelength is defined by the Shannon entropy $H$ over the nodes’ visit rates [42], $H\left(P\right)=-\sum_{u\in V}p_{u}\log_{2}p_{u}$, where $p_{u}$ is node $u$’s visit rate and $P=\left\{p_{u}\,|\,u\in V\right\}$ is the set of node visit rates. In undirected graphs, we compute visit rates directly as $p_{u}=s_{u}/\sum_{v\in V}s_{v}$, where $s_{u}=\sum_{v\in V}\delta\left(u,v\right)$ is node $u$’s strength. In directed graphs, we compute the visit rates numerically with smart teleportation [43] and a power iteration. When we partition the nodes into modules, the codelength becomes a weighted average of the modules’ entropies and the entropy at the so-called index level for switching between modules. Figure 1 illustrates the coding principle behind the map equation using Huffman codes [44]; note, however, that these codewords are only for illustration and we only care about their expected length in the limit to evaluate the map equation.
@@
Here, $q=\sum_{\mathsf{m}}q_{\mathsf{m}}$ is the random walker’s module entry rate, $q_{\mathsf{m}}$ is module $\mathsf{m}$’s entry rate, and $Q=\left\{q_{\mathsf{m}}\,\middle|\,\mathsf{m}\in\mathsf{M}\right\}$ is the set of module entry rates; $p_{\mathsf{m}}=\mathsf{m}_{\text{exit}}+\sum_{u\in\mathsf{m}}p_{u}$ is the rate at which the random walker moves in module $\mathsf{m}$, including the module exit rate $\mathsf{m}_{\text{exit}}$, and $P_{\mathsf{m}}=\left\{\mathsf{m}_{\text{exit}}\right\}\cup\left\{p_{u}\,\middle|\,u\in\mathsf{m}\right\}$ is the set of node visit and exit rates for module $\mathsf{m}$. We can rewrite the map equation as [18]
$L\left(\mathsf{M}\right)=q\log_{2}q-2\sum_{\mathsf{m}\in\mathsf{M}}q_{\mathsf{m}}\log_{2}q_{\mathsf{m}}-\sum_{u\in V}p_{u}\log_{2}p_{u}+\sum_{\mathsf{m}\in\mathsf{M}}p_{\mathsf{m}}\log_{2}p_{\mathsf{m}}.$ (2)
The map equation framework has been extended for overlapping communities through state-space expansions with higher-order network models [45, 46], avoid over-partitioning in sparse networks using a Bayesian regularisation approach [47], and to deal with sparse constrained structures [48]. Moreover, the map equation framework can incorporate node features through an extension [49] or by preprocessing data [50]. Detecting communities relies on Infomap [14], a greedy stochastic search algorithm that optimises the map equation. However, each of the above extensions requires preprocessing the input data, adjusting the loss function, or adapting the search algorithm. In contrast, adapting the map equation as a loss function for optimisation with gradient descent does not require any custom algorithm, thus enabling flexible experimentation with variations, scalability to GPU clusters, and incorporating it into other loss functions.
@@ 4 The map equation goes neural
We set out to detect communities by optimising the map equation with GNNs through gradient descent, which essentially means learning coarse-graining node representations in the form of communities (Figure 2). While the standard map equation considers hard clusters where each node is assigned to exactly one module, we introduce a soft cluster assignment matrix $\mathbf{S}_{n\times s}$ to make the map equation differentiable and enable overlapping clusters. We optimise $\mathbf{S}=\operatorname{softmax}\left(\operatorname{GNN}_{\theta}\left(\mathbf{A},\mathbf{X}\right)\right)$ indirectly by optimising the GNN’s parameters $\theta$, that is weights, with respect to the codelength $L$. Here, $\mathbf{A}_{n\times n}$ is the graph’s adjacency matrix, $\mathbf{X}_{n\times d}$ is the node features matrix, $n=\left|V\right|$ is the number of nodes, $s$ is the maximum allowed number of clusters, and $d$ is the node feature dimension.
@@
Without loss of generality, we assume directed networks. We denote the weight on the link from node $i$ to $j$ as $w_{ij}=\delta\left(i,j\right)$ and the graph’s total weight as $w_{\text{tot}}=\sum_{i\in V}\sum_{j\in V}w_{ij}$. Let $\mathbf{T}_{n\times n}$ be the random walker’s transition matrix and $\mathbf{d}^{\text{in}}$ be the vector of weighted node in-degrees with
@@
$q=1-\operatorname{tr}\left(\mathbf{C}\right),\qquad\mathbf{q}_{\mathsf{m}}=\mathbf{C}\mathbf{1}_{s}-\operatorname{diag}\left(\mathbf{C}\right),\qquad\mathbf{p}_{\mathsf{m}}=\mathbf{q}_{\mathsf{m}}+\mathbf{1}_{s}^{\top}\mathbf{C},$
@@
$L\left(\mathbf{A},\mathbf{S}\right)=q\log_{2}q-2\left(\mathbf{q}_{\mathsf{m}}\log_{2}\mathbf{q}_{\mathsf{m}}\right)\mathbf{1}_{s}+\left(\mathbf{p}\log_{2}\mathbf{p}\right)\mathbf{1}_{n}+\left(\mathbf{p}_{\mathsf{m}}\log_{2}\mathbf{p}_{\mathsf{m}}\right)\mathbf{1}_{s},$ (3)
@@
The most expensive calculation is the pooling operation $\mathbf{C}=\mathbf{S}^{\top}\mathbf{F}\mathbf{S}$ which depends on the network’s sparsity. When $s\ll n$ and the number of edges $m=\mathcal{O}\left(n\right)$, the complexity of Neuromap is linear in $n$. When the network is dense, $m=\mathcal{O}\left(n^{2}\right)$, or the maximum number of clusters approaches the number of nodes $s\approx n$, we approach quadratic complexity. Therefore, we recommend keeping $s\ll n$ for scalability.
@@ 5 Experimental evaluation
We evaluate Neuromap on synthetic and real-world networks with different neural network architectures: a 2-layer graph convolutional network (GCN) [6], a 2-layer graph isomorphism network (GIN) [52], and a 2-layer SAGE network [53]. To investigate whether GNNs are required for clustering, we also include a fully connected linear layer and a 2-layer MLP. In all cases, we use the models provided by PyTorch Geometric with SELU activation [54]. We use the Adam optimiser [55], apply batch normalisation, and for comparability between different methods, set the learning rate for the linear layer to $10^{-1}$, for MLP to $10^{-2}$, and for GCN, GIN, and SAGE to $10^{-3}$. We train all models for up to $10{,}000$ epochs with a patience of $100$ epochs and dropout probability $0.5$. Because the datasets contain hard clusters, we convert the resulting communities to hard clusters, assigning each node to that cluster where it has its strongest membership. As baselines, we use Infomap [14] and five recent approaches for unsupervised graph clustering with GNNs: DMoN [34], NOCD [36], DiffPool [7], MinCut [8], and Ortho [34]. We base our implementation [^1] https://github.com/chrisbloecker/neuromap on PyTorch [56] and PyTorch Geometric [57] and run our experiments on a workstation with an Intel i9-11900K @ 3.50GHz CPU, 32 GB of RAM, and a GeForce RTX 3090 with 24 GB of memory.
@@ 5.1 Synthetic networks with planted communties
We generate both directed and undirected Lancichinetti-Fortunato-Radicchi (LFR) benchmark networks with planted ground-truth communities [58] with $n=1000$ nodes, average node degree $k\in\left\{\ln{n},2\ln{n}\right\}$, maximum node degree $k_{\operatorname{max}}=2\sqrt{n}$, both rounded to the nearest integer, and mixing parameter $\mu$ between $0.1$ and $0.8$ with a step size of $0.1$. We set the power-law exponents for the node degree distribution to $\tau_{1}=2$, and for the community size distribution to $\tau_{2}=1$. For each combination of parameters, we generate $10$ LFR networks using the implementation provided by the authors, [^2] https://sites.google.com/site/andrealancichinetti/benchmarks resulting in a total of 320 networks. For each parameter combination, there are 10 LFR networks; for each of these LFR networks, we run each model 10 times, measuring its performance as the average adjusted mutual information (AMI) [59] against the ground truth, and plot the average of those AMI values over the 10 networks per parameter combination. To verify that the number of communities is inferred from the data, we set the maximum number of communities to $s=n$. Since LFR networks do not have node features, we use the adjacency matrix as node features.
@@
We compare Neuromap against recent deep-learning-based community detection methods on the same networks by swapping out the loss function while keeping everything else the same, with the exception of using weight decay for NOCD as per the original paper [36]. For DiffPool, Mincut, Ortho, and DMoN, we use the implementation from PyTorch Geometric, for NOCD, we use the implementation provided by the authors. [^3] https://github.com/shchur/overlapping-community-detection.git Figure 4 shows the results for SAGE; in Appendix A, we also include results for the remaining architectures. Neuromap outperforms all baselines with all architectures, except for NOCD on directed networks with GCN. Different from previous works, we have not limited the maximum number of communities, which allows us to analyse the methods’ overfitting behaviour: While Neuromap reports close to the ground truth number of communities, the remaining methods overfit the networks’ structure and report considerably more communities, which is especially pronounced for Ortho and DMoN as well as for NOCD when using SAGE on directed networks (note the logarithmic scale). MinCut fails to identify meaningful communities on directed networks for mixing values $\mu>0.3$. NOCD performs best with the GCN architecture, which was also used in the original paper [36]. Neuromap performs best with SAGE in our experiments.

### v4

@@ Abstract
Community detection is an essential tool for unsupervised data exploration and revealing the organisational structure of networked systems. With a long history in network science, community detection typically relies on objective functions, optimised with custom-tailored search algorithms, but often without leveraging recent advances in deep learning. Recently, first works have started incorporating such objectives into loss functions for deep graph clustering and pooling. We consider the map equation, a popular information-theoretic objective function for unsupervised community detection, and express it in differentiable tensor form for optimisation through gradient descent. Our formulation turns the map equation compatible with any neural network architecture, enables end-to-end learning, incorporates node features, and chooses the optimal number of clusters automatically, all without requiring explicit regularisation. Applied to unsupervised graph clustering tasks, we achieve competitive performance against state-of-the-art deep graph clustering baselines in synthetic and real-world datasets.
@@ 1 Introduction
Many real-world networked systems are organised in communities: groups of nodes that are more similar to each other than to the rest. Communities provide insights into network structure at the mesoscale, revealing sub-systems by analysing link patterns. Motivated by different research questions, several characterisations of what constitutes “good” communities have been proposed [1, 2], however, neither of them is fundamentally more correct than any other. Moreover, no single community-detection method outperforms all others on any given network [3], motivating the ongoing efforts of research on community detection. Typically, community-detection approaches formulate an objective function that calculates a quality score for a given partition of the network’s nodes into communities. Finding the best partition is an NP-hard search problem and often involves custom heuristic algorithms that attempt to minimise their objective function [4, 5, 6].
@@
Here, we combine the benefits of traditional community-detection approaches and deep learning and consider the map equation, an information-theoretic objective function for community detection [16]. By adapting the map equation for soft cluster assignments and implementing it in differentiable tensor form, we enable end-to-end optimisation of the map equation as a loss function with gradient descent and GNNs. In analogy to the map equation’s stochastic optimisation algorithm Infomap [6], we call our approach Neuromap and evaluate it against Infomap and several recent GNN-based graph clustering methods. Applied to synthetic and real-world networks, Neuromap demonstrates competitive performance against recent deep graph clustering baselines.
@@
    We adapt the map equation as a differentiable loss function for end-to-end deep graph clustering and propose Neuromap, a deep-learning-based alternative to the popular Infomap algorithm for unsupervised community detection with the map equation. Neuromap is compatible with any neural network architecture, detects overlapping communities, leverages node features for improved performance on real-world networks, and, by following the minimum description length principle, does not require explicit regularisation.
@@
    We extensively evaluate Neuromap on hundreds of synthetic and ten real datasets against recent baselines paired with various neural network architectures. Neuromap outperforms the baseline on the synthetic networks in most settings and is amongst the best performers in seven out of ten real datasets.
@@
    By choosing a higher maximum number of clusters than previous works, we show empirically that recent baselines tend to overfit and report considerably more than the ground-truth number of communities. Moreover, we find that choosing a small maximum number of communities is often detrimental to graph clustering performance.
@@ 3 Background: the map equation
Let $G=\left(V,E\right)$ be a graph with nodes $V$, links $E$, and let $w_{uv}\in\mathbb{R}^{+}_{0}$ denote the non-negative link weight on the link from node $u$ to $v$. When all nodes are assigned to the same module, the codelength is defined by the Shannon entropy $H$ over the nodes’ visit rates [44], $H\left(P\right)=-\sum_{u\in V}p_{u}\log_{2}p_{u}$, where $p_{u}$ is node $u$’s visit rate and $P=\left\{p_{u}\,|\,u\in V\right\}$ is the set of node visit rates. In undirected graphs, we compute visit rates directly as $p_{u}=s_{u}/\sum_{v\in V}s_{v}$, where $s_{u}=\sum_{v\in V}w_{uv}$ is node $u$’s strength. In directed graphs, we compute the visit rates numerically with smart teleportation [45] and a power iteration. When we partition the nodes into modules, the codelength becomes a weighted average of the modules’ entropies and the entropy at the so-called index level for switching between modules. Figure 1 illustrates the coding principle behind the map equation using Huffman codes [46]; note, however, that these codewords are only for illustration and we only care about their expected length in the limit to evaluate the map equation.
@@
Here, $q=\sum_{\mathsf{m}}q_{\mathsf{m}}$ is the random walker’s module entry rate, $q_{\mathsf{m}}=\sum_{u\notin\mathsf{m}}\sum_{v\in\mathsf{m}}p_{u}t_{uv}$ is module $\mathsf{m}$’s entry rate, and $Q=\left\{q_{\mathsf{m}}/q\mid\mathsf{m}\in\mathsf{M}\right\}$ is the set of normalised module entry rates; $p_{\mathsf{m}}=\mathsf{m}_{\text{exit}}+\sum_{u\in\mathsf{m}}p_{u}$ is the rate at which the random walker moves in module $\mathsf{m}$, including the module exit rate $\mathsf{m}_{\text{exit}}=\sum_{u\in\mathsf{m}}\sum_{v\notin\mathsf{m}}p_{u}t_{uv}$, and $P_{\mathsf{m}}=\left\{\mathsf{m}_{\text{exit}}/p_{\mathsf{m}}\right\}\cup\left\{p_{u}/p_{\mathsf{m}}\mid u\in\mathsf{m}\right\}$ is the set of normalised node visit and exit rates for module $\mathsf{m}$. The random walker’s transition probability from node $u$ to $v$ is $t_{uv}=w_{uv}/\sum_{v\in V}w_{uv}$. We can rewrite the map equation as (see Appendix A)
$L\left(\mathsf{M}\right)=q\log_{2}q-\!\sum_{\mathsf{m}\in\mathsf{M}}q_{\mathsf{m}}\log_{2}q_{\mathsf{m}}-\!\sum_{\mathsf{m}\in\mathsf{M}}\mathsf{m}_{\text{exit}}\log_{2}\mathsf{m}_{\text{exit}}-\!\sum_{u\in V}p_{u}\log_{2}p_{u}+\!\sum_{\mathsf{m}\in\mathsf{M}}p_{\mathsf{m}}\log_{2}p_{\mathsf{m}}.$ (2)
The map equation framework has been extended for overlapping communities through state-space expansions with higher-order network models [47, 48], avoiding over-partitioning in sparse networks using a Bayesian regularisation approach [49], and to deal with sparse constrained structures [50]. Moreover, the map equation framework can incorporate node features through an extension [51] or by preprocessing data [52]. Detecting communities relies on Infomap [6], a greedy stochastic search algorithm that optimises the map equation. However, each of the above extensions requires preprocessing the input data, adjusting the loss function, or adapting the search algorithm. In contrast, adapting the map equation as a loss function for optimisation with gradient descent does not require any custom algorithm, thus enabling flexible experimentation with variations, scalability to GPU clusters, and incorporating it into other loss functions.
@@ 4 The map equation goes neural
We set out to detect communities by optimising the map equation with GNNs through gradient descent, which essentially means learning coarse-graining node representations in the form of communities (Figure 2). While the standard map equation considers hard clusters where each node is assigned to exactly one module, we introduce a soft cluster assignment matrix $\mathbf{S}_{n\times s}$ to make the map equation differentiable and enable overlapping clusters. We optimise $\mathbf{S}=\operatorname{softmax}\left(\operatorname{GNN}_{\theta}\left(\mathbf{A},\mathbf{X}\right)\right)$ indirectly by optimising the GNN’s parameters $\theta$, that is, its weights, with respect to the codelength $L$. Here, $\mathbf{A}_{n\times n}$ is the graph’s adjacency matrix, $\mathbf{X}_{n\times d}$ is the node features matrix, $n=\left|V\right|$ is the number of nodes, $s$ is the maximum allowed number of clusters, and $d$ is the node feature dimension.
@@
Without loss of generality, we assume directed networks. We denote the graph’s total weight as $w_{\text{tot}}=\sum_{i\in V}\sum_{j\in V}w_{ij}$. Let $\mathbf{T}_{n\times n}$ be the random walker’s transition matrix and $\mathbf{d}^{\text{in}}$ be the vector of weighted node in-degrees with
@@
$q=1-\operatorname{tr}\left(\mathbf{C}\right)\quad\mathbf{q}_{\mathsf{m}}=\mathbf{C}\mathbf{1}_{s}-\operatorname{diag}\left(\mathbf{C}\right)\quad\mathbf{m}_{\text{exit}}=(\mathbf{1}_{s}^{\top}\mathbf{C})^{\top}\!-\operatorname{diag}\left(\mathbf{C}\right)\quad\mathbf{p}_{\mathsf{m}}=\mathbf{q}_{\mathsf{m}}+\mathbf{1}_{s}^{\top}\mathbf{C}$
@@
$L\left(\mathbf{A},\mathbf{S}\right)=q\log_{2}q-\left(\mathbf{q}_{\mathsf{m}}\log_{2}\mathbf{q}_{\mathsf{m}}\right)\mathbf{1}_{s}-\left(\mathbf{m}_{\text{exit}}\log_{2}\mathbf{m}_{\text{exit}}\right)\mathbf{1}_{s}-\left(\mathbf{p}\log_{2}\mathbf{p}\right)\mathbf{1}_{n}+\left(\mathbf{p}_{\mathsf{m}}\log_{2}\mathbf{p}_{\mathsf{m}}\right)\mathbf{1}_{s}$ (3)
@@
The most expensive calculation is the pooling operation $\mathbf{C}=\mathbf{S}^{\top}\mathbf{F}\mathbf{S}$ which depends on the network’s density. When $s\ll n$ and the number of edges is $m=\mathcal{O}\left(n\right)$, the complexity of Neuromap is linear in $n$. When the network is dense, $m=\mathcal{O}\left(n^{2}\right)$, or the maximum number of clusters approaches the number of nodes $s\approx n$, we approach quadratic complexity. Therefore, we recommend keeping $s\ll n$ for scalability.
@@ 5 Experimental evaluation
We evaluate Neuromap on synthetic and real-world networks with different neural network architectures: a 2-layer graph convolutional network (GCN) [9], a 2-layer graph isomorphism network (GIN) [54], and a 2-layer SAGE network [55]. To investigate whether GNNs are required for clustering, we also include a fully connected linear layer and a 2-layer MLP. We include a learnable temperature parameter for the softmax operation, which we found speeds up convergence. In all cases, we use the models provided by PyTorch Geometric with SELU activation [56]. Because the specifics between architectures differ, such as message-passing details and aggregation functions, they may be interpreted as using different search algorithms which return different communities. We use the Adam optimiser [57], apply batch normalisation, and for comparability between different methods, set the learning rate for the linear layer to $10^{-1}$, for MLP to $10^{-2}$, and for GCN, GIN, and SAGE to $10^{-3}$. We train all models for up to $10{,}000$ epochs with a patience of $100$ epochs and dropout probability $0.5$. Because the datasets contain hard clusters, we convert the resulting communities to hard clusters, assigning each node to that cluster where it has its strongest membership. As baselines, we use Infomap [6] and five recent approaches for unsupervised graph clustering with GNNs: DMoN [36], NOCD [38], DiffPool [10], MinCut [11], and Ortho [36]. We base our implementation [^1] https://github.com/chrisbloecker/neuromap on PyTorch [58] and PyTorch Geometric [59] and ran our experiments on a workstation with an Intel i9-11900K @ 3.50GHz CPU, 32 GB of RAM, and a GeForce RTX 3090 with 24 GB of memory.
@@ 5.1 Synthetic networks with planted communties
We generate directed and undirected Lancichinetti-Fortunato-Radicchi (LFR) benchmark networks with planted ground-truth communities [60] with $n=1000$ nodes, average node degree $k\in\left\{\ln{n},2\ln{n}\right\}$, maximum node degree $k_{\operatorname{max}}=2\sqrt{n}$, both rounded to the nearest integer, and mixing parameter $\mu$ between $0.1$ and $0.8$ with a step size of $0.1$. We set the power-law exponents for the node degree distribution to $\tau_{1}=2$, and for the community size distribution to $\tau_{2}=1$. For each combination of parameters, we generate $10$ LFR networks using the implementation provided by the authors, [^2] https://sites.google.com/site/andrealancichinetti/benchmarks resulting in a total of 320 networks. For each parameter combination, there are 10 LFR networks; for each of these LFR networks, we run each model 10 times, measuring its performance as the average adjusted mutual information (AMI) [61] against the ground truth, and plot the average of those AMI values over the 10 networks per parameter combination. To verify that the number of communities is inferred from the data, we set the maximum number of communities to $s=n$. Since LFR networks do not have node features, we use the adjacency matrix as node features.
@@
We compare Neuromap against recent deep-learning-based community detection methods on the same networks by swapping out the loss function while keeping everything else the same, with the exception of using weight decay for NOCD as per the original paper [38]. For DiffPool, Mincut, Ortho, and DMoN, we use the implementation from PyTorch Geometric, for NOCD, we use the implementation provided by the authors. [^3] https://github.com/shchur/overlapping-community-detection Figure 4 shows the results for SAGE; in Appendix B, we also include results for the remaining architectures. Neuromap outperforms the baselines across all architectures, except for NOCD which performs better than Neuromap with GIN and with GCN on directed networks. Different from previous works, we have not limited the maximum number of communities, which allows us to analyse the methods’ overfitting behaviour: While Neuromap reports close to the ground-truth number of communities, the remaining methods often overfit the networks’ structure and report considerably more communities (note the logarithmic scale). MinCut fails to identify meaningful communities on directed networks for mixing values $\mu>0.3$. NOCD performs best with the GCN architecture, which was also used in the original paper [38]. Neuromap performs best with SAGE in our experiments.